# Note on the Humbert function $\Psi_{1}$

### v1

@@ 3. Proof of the main results
Recall the series representation (1.2). We have $V_{1}(x,y)=\mathcal{O}(1)$ since $\frac{y}{1-x}$ is bounded and then $V_{1}(x,y)$ is absolutely convergent. Hence the main contribution of $\Psi_{1}$ comes from $V_{2}(x,y)$. Now divide $V_{2}(x,y)$ into two parts:
@@
For $10\leqslant n\leqslant N-$ 1, recall the exponential expansion of ${}_{2}F_{2}$ [6, equation (5.8)]
@@
as $y\to+\infty$, where the coefficients $c_{k,n}$ are given by [10, equation (12)]
@@
$V_{2}^{(L)}(x,y)=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-b-c^{\prime}}\mathrm{e}^{y}\sum_{n=0}^{N-1}\frac{\left(b\right)_{n}}{n!}\left(1-x\right)^{-n}\sum_{k=0}^{N-1}c_{k,n}^{*}y^{-k}+\mathcal{O}\left(y^{\Re(a+b-c-d)-N}\mathrm{e}^{y}\right),$
@@
$\left(1-x\right)^{-b}V_{2}(x,y)$ $=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{p-N}\mathrm{e}^{y}\right)$
$=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{M-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{\Re(a-b-c^{\prime})-M}\mathrm{e}^{y}\right)+\mathcal{O}\left(y^{p-N}\mathrm{e}^{y}\right)$
$=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{M-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{\Re(a-b-c^{\prime})-M}\mathrm{e}^{y}\right),$
@@
Recall the asymptotic expansion of ${}_{2}F_{2}(y)$ [6, equation (5.8)]: for $0\leqslant n\leqslant N$,
@@
where $a_{3}=1-b,\,b_{3}=1-d$, $S_{n}(-y)$ and $T_{n}(-y)$ are defined in Theorem 2.4, and the coefficients $c_{k,n}$ are given by (3.3). For $n\geqslant N+1$, recall the expansion (2.2). Use of (2.1) and the definition of $F(x,y)$ then gives
@@
applying the transformation [7, p. 100, equation (1)]
@@
$t_{k}^{-}(x,y)$ $={}_{4}F_{5}\left[\begin{matrix}a_{1}+k,a_{2}+k,a_{3}+k,a-b\\ b_{1}+k,b_{2}+k,b_{3}+k,1-b,c-b\end{matrix};\frac{y}{1-x}\right]\left(\frac{y}{x-1}\right)^{k}.$
@@ 4. Concluding remarks
We have established the complete asymptotic expansions of the Humbert function $\Psi_{1}$ for two large arguments. Our derivation is based on the estimates of ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ for large $z$ (see Theorems 2.42.6), but the exponential expansions are not included.
We conjecture that the more accurate asymptotic expansion (3.6) is valid for $n\in\mathbb{Z}$ with the remainder terms multiplied by $n^{\lambda}$ for some $\lambda>0$. If this is the case, we may deal with some revelant problems:
  • (i)
    Asymptotics of $\Psi_{1}[a,b;c,c^{\prime};x,y]$ for large $y$: This problem was partially solved in [4, Section 3.1]. We guess that the series representations (1.2) and [3, equation (3.4)] offer the complete asymptotic expansions.
  • (ii)
    Asymptotics of $\Phi_{3}[b;c;w,z]$ for large $z$ : One may deduce the asymptotic behavior via the series representation [2, equation (1.13)]
    $\Phi_{3}[b;c;w,z]=\sum_{m=0}^{\infty}\frac{\left(b\right)_{m}}{\left(c\right)_{m}}{}_{1}F_{1}\left[\begin{matrix}-m\\ 1-b-m\end{matrix};\frac{z}{w}\right]\frac{w^{m}}{m!}.$
  • (iii)
    Asymptotics of $\Phi_{1},\Phi_{2},\Xi_{1}$ for large $y$: See the series representations shown in [5, p. 76].
Braaksma’s work [1] seems helpful to examine this conjecture.
Finally, we are not aware right now of any application of our Theorems 1.2 and 1.3.
@@ Acknowledgements
The author thanks Prof. Nico M. Temme for bringing Braaksma’s work to his attention.

### v3

@@ 3.  Proof of the main results
Recall the series representation (1.3). We have $V_{1}(x,y)=\mathcal{O}(1)$ since $\frac{y}{1-x}$ is bounded. Hence the main contribution of $\Psi_{1}$ comes from $V_{2}(x,y)$. Now divide $V_{2}(x,y)$ into two parts:
@@
For $10\leqslant n\leqslant N-$ 1, recall the exponential expansion of ${}_{2}F_{2}$ [7, Equation (5.8)]
@@
as $y\to+\infty$, where the coefficients $c_{k,n}$ are given by [11, Equation (12)]
@@
$V_{2}^{(L)}(x,y)=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-b-c^{\prime}}\mathrm{e}^{y}\sum_{n=0}^{N-1}\frac{\left(b\right)_{n}}{n!}\left(1-x\right)^{-n}\sum_{k=0}^{N-1}c_{k,n}^{*}y^{-k}+\mathcal{O}\left(y^{\Re(a-b-c^{\prime})-N}\mathrm{e}^{y}\right),$
@@
$\left(1-x\right)^{-b}V_{2}(x,y)$ $=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{\Re(a-2b-c^{\prime})-N}\mathrm{e}^{y}\right)+\mathcal{O}\left(y^{p-\Re(b)-N}\mathrm{e}^{y}\right)$
$=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{M-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{\Re(a-2b-c^{\prime})-M}\mathrm{e}^{y}\right)+\mathcal{O}\left(y^{p-\Re(b)-N}\mathrm{e}^{y}\right)$
$=\frac{\Gamma(c^{\prime})}{\Gamma(a-b)}y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{M-1}a_{k}(x,y)y^{-k}+\mathcal{O}\left(y^{\Re(a-2b-c^{\prime})-M}\mathrm{e}^{y}\right),$
@@
Recall the asymptotic expansion of ${}_{2}F_{2}(y)$ [7, Equation (5.8)]: for $0\leqslant n\leqslant N$,
@@
where $a_{3}=1-b,\,b_{3}=1-d$, $S_{n}(-y)$ and $T_{n}(-y)$ are defined in Theorem 2.4, and the coefficients $c_{k,n}$ are given by (3.3). For $n\geqslant N+1$, recall the expansion (2.2). Using (2.1) and the definition of $F(x,y)$ then gives
@@
applying the transformation [8, p. 100, Equation (1)]
@@
$t_{k}^{-}(x,y)$ $={}_{5}F_{5}\left[\begin{matrix}a_{1}+k,a_{2}+k,a_{3}+k,a-b,1\\ b_{1}+k,b_{2}+k,b_{3}+k,1-b,c-b\end{matrix};\frac{y}{1-x}\right]\left(\frac{y}{x-1}\right)^{k}.$
@@ 4.  Concluding remarks
We have established the complete asymptotic expansions of the Humbert function $\Psi_{1}$ for two large arguments. Our derivation is based on the estimates of ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ for large $z$ (see Theorems 2.42.6), but the exponential expansions are not included. We conjecture that the more accurate expansion (3.6) is valid for $n\in\mathbb{Z}$ with remainders multiplied by $n^{\lambda}$ for some $\lambda>0$. Braaksma’s work [1] seems helpful to examine this conjecture.
Regarding asymptotics of $\Psi_{1}$ for large $y$, it was partially studied in [5, Section 3.1]. The series representations (1.3) and [4, Equation (3.4)] appear to provide complete asymptotic expansions for large $y$ with unrestricted $\arg(x)$ and $\arg(y)$. We will determine appropriate methods to verify this. Additionally, the asymptotics of other Humbert functions will be explored in future studies.
Finally, we want to make further remarks on our Theorems 1.2 and 1.3. Note that $\Psi_{1}$ has a Kummer-type transformation [3, Equation (2.54)]
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\left(1-x\right)^{-a}\Psi_{1}\left[a,c-b;c,c^{\prime};\frac{x}{x-1},\frac{y}{1-x}\right],$ (4.1)
which elucidates that the singularity of $\Psi_{1}[x,y]$ as $x\to 1$ is equivalent to its asymptotics under the condition (1.5). As an example, one may get from (4.1) and Theorem 1.2 that for fixed $y>0$,
$\Psi_{1}\left[1,\frac{1}{2};\frac{3}{2},\frac{1}{2};x,y\right]\sim\frac{\sqrt{\pi}}{2}y^{-\frac{1}{2}}\left(1-x\right)^{\frac{1}{2}}\mathrm{e}^{\frac{y}{1-x}},\quad x\to 1^{-},$ (4.2)
which is confirmed by [4, Equation (3.4)] and can be numerically verified by the integral [2, p. 115, Equation (38)]
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\frac{\Gamma(c)}{\Gamma(b)\Gamma(c-b)}\int_{0}^{1}u^{b-1}\left(1-u\right)^{c-b-1}\left(1-ux\right)^{-a}{}_{1}F_{1}\left[\begin{matrix}a\\ c^{\prime}\end{matrix};\frac{y}{1-ux}\right]\mathrm{d}u,$
where $\Re(c)>\Re(b)>0$ and $(x,y)\in\mathbb{D}_{\Psi_{1}}$ (see (1.1)).
@@ Acknowledgements
We would like to thank Malte Henkel for suggesting the inclusion of the example (4.2), and Nico Temme for bringing Braaksma’s work to our attention in a personal discussion with the first author.